We dissolve 2.77 g of an unknown acid, HA, in enough water to produce 25.0 mL of solution. The pH of this solution of HA(aq) is 1.33. We titrate this solution with a 0.250 M solution of NaOH. It takes 44.5 mL of the NaOH solution to reach the equivalence point.

(a) What is the molar mass of HA?

(b) What is the pKa value of HA(aq)?

(c) What is the pH at the equivalence point?

(d) What is the pH one third along the titration when you still have twice as much HA(aq) as A-(aq)?
Answer #1

- b) 0.0ll1a5 mol 25.0x103 L 0.445 m molarity of HA = PH = -4096,0) = -loofkac) -) Skae = 10² 10-PH =10-33 2 Kac ka C = 2.30N.B.- If you have any further queries, please feel free to ask in the comments section.